OpenStudy (anonymous):

Write the equation of the line with a slope of -5 and a y-intercept of (0,3)

OpenStudy (igreen):

Put it in point-slope form first. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf y_1\) is the y-value of the point, \(\sf x_1\) is the x-value of the point, and \(\sf m\) is the slope.

OpenStudy (anonymous):

so 3-0=-5(3-0)

OpenStudy (igreen):

No, we don't do anything to 'x' or 'y'

OpenStudy (anonymous):

so

OpenStudy (igreen):

(0, 3) x1 y1

OpenStudy (igreen):

m = -5

OpenStudy (igreen):

You got the slope right. \(\sf y - y_1 = -5(x - x_1)\)

OpenStudy (anonymous):

okay. so y-3=-5(x-0)

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

and that is the equation?

OpenStudy (igreen):

Yes, do you have choices? It might want it in slope-intercept form.

OpenStudy (anonymous):

nah. I don't think it does because it doesn't say anything about slope intercept